Certain corresponding values of x and log10x are given below:

x: 300 304 305 307

log10 x: 2.4771 2.4829 2.4843 2.4871

Find log10 (310) by Lagrange’s formula.


1
Expert's answer
2021-01-11T18:03:35-0500

Here the intervals are unequal.


"\begin{matrix}
 x_0=300 & y_0=2.4771 \\
\\
 x_1=304 & y_1=2.4829 \\
\\
 x_2=305 & y_2=2.4843 \\
\\
 x_3=307 & y_3=2.4871 \\
\end{matrix}"


By Lagrange’s interpolation formula we have

"y=f(x)=\dfrac{(x-x_1)(x-x_2)(x-x_3)}{(x_0-x_1)(x_0-x_2)(x_0-x_3)}\times y_0"

"+\dfrac{(x-x_0)(x-x_2)(x-x_3)}{(x_1-x_0)(x_1-x_2)(x_1-x_3)}\times y_1"

"+\dfrac{(x-x_0)(x-x_1)(x-x_3)}{(x_2-x_0)(x_2-x_1)(x_2-x_3)}\times y_2"

"+\dfrac{(x-x_0)(x-x_1)(x-x_2)}{(x_3-x_0)(x_3-x_1)(x_3-x_2)}\times y_3"



Put "x=310"

"f(310)=\dfrac{(310-304)(310-305)(310-307)}{(300-304)(300-305)(300-307)}\times 2.4771"

"+\dfrac{(310-300)(310-305)(310-307)}{(304-300)(304-305)(304-307)}\times 2.4829"




"+\dfrac{(310-300)(310-304)(310-307)}{(305-300)(305-304)(305-307)}\times 2.4843"

"+\dfrac{(310-300)(310-304)(310-305)}{(307-300)(307-304)(307-305)}\times 2.4871"

"=2.4914"

"\log_{10}(310)=2.4914"